OpenStudy (anonymous):

multiply (4sqrt(2)+sqrt(6))(8sqrt(2)-9sqrt(6))

OpenStudy (anonymous):

\[(4\sqrt{2}+\sqrt{6})(8\sqrt{2}-9\sqrt{6})\]

OpenStudy (anonymous):

\[(4*8*2)-(36\sqrt{12})+(8\sqrt{12})-(9*6)\]

OpenStudy (anonymous):

\[64-(72\sqrt{3})+(16\sqrt{3})-54\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[(4\sqrt{2}+\sqrt{6})(8\sqrt{2}-9\sqrt{6})\]

OpenStudy (anonymous):

\[4\sqrt2\times 8\sqrt2=4\times 8\times 2=64\]

OpenStudy (anonymous):

oh i see you did most of it already

OpenStudy (anonymous):

yesm :) just lost on that one part

OpenStudy (anonymous):

actually you did all of it know when to stop!

OpenStudy (anonymous):

\[10-56\sqrt{3}\] finished

OpenStudy (anonymous):

roger

OpenStudy (anonymous):

that negative signage

OpenStudy (anonymous):

what about it?

OpenStudy (anonymous):

is what confused me -72+16 is all it was

OpenStudy (anonymous):

oh ok